# From Image to UML: First Results of Image-Based UML Diagram Generation using LLMs

### v1

@@ Abstract
In software engineering processes, systems are first specified using a modeling language such as UML. These initial designs are often collaboratively created, many times in meetings where different domain experts use whiteboards, paper or other types of quick supports to create drawings and blueprints that then will need to be formalized. These proper, machine-readable, models are key to ensure models can be part of automated processes (e.g. input of a low-code generation pipeline, a model-based testing system, …). But going form hand-drawn diagrams to actual models is a time-consuming process that sometimes ends up with such drawings just added as informal images to the software documentation, reducing their value a lot. To avoid this tedious task, we explore the usage of Large Language Models (LLM) to generate the formal representation of (UML) models from a given drawing. More specifically, we have evaluated the capabilities of different LLMs to convert images of UML class diagrams into the actual models represented in the images. While the results are good enough to use such an approach as part of a model-driven engineering pipeline we also highlight some of their current limitations and the need to keep the human in the loop to overcome those limitations.
Keywords Large Language Model  $\cdot$ UML Diagram  $\cdot$ Software Models  $\cdot$ Low-code
@@ 1 Introduction
The continuous progress and rise of Large Language Models (LLM) has led to their integration into various tasks in various domains, such as medical advice consultation in healthcare, writing or reading assistance in education, legal interpretation and reasoning in Law or financial reasoning [1]. Computer science is not an exception, with various LLMs being used in software development as programming assistants [2, 3], but also in software modeling as a model creation tool from natural language [4, 5], showing promising results on both fronts.
@@
Beyond textual input, LLMs such as ChatGPT4 or Google’s Gemini additionally provide support for image input, also described as visual LLMs. Generally, OpenAI’s GPT-4V [^1] https://openai.com/research/gpt-4v-system-card and Google’s Gemini (Pro/Ultra) [^2] https://deepmind.google/technologies/gemini Vision are seen as the best publicly available multimodal LLM. Unfortunately, these are either hidden behind a paywall or the free version is only accessible via a provided platform, but not accessible as a stand alone service for free, such as Gemini Pro being freely accessible only on Google’s chat interface. CogVLM [6] acts as an open-source alternative, providing not only its source code but also a free chat interface to interact with the model. These multi-modal LLMs further broadened the range of possible applications, such as tools like Design2Code [7] or Make-real [^3] https://github.com/tldraw/make-real that allow for the generation of HTML, CSS and JS code from either screenshots of web pages or mock-ups. In the domain of software modeling, we believe that these image reading capabilities could also be leveraged to facilitate and accelerate the modeling process itself.
@@
Indeed, the adoption of modeling tools is constantly debated [8], with one of the biggest hurdles being the usability of such tools [9]. Take the example of UML class diagrams that are used to represent the structure of software. These are usually collaboratively sketched on a whiteboard in an attempt to crystallize the initial design idea of an application. Yet, once drawn, there still is the need to transform the drawing into a more polished form, for example for documentation purposes. Additionally, one might also want a computer readable format to use the UML class diagram for code generation. The same goes for migrating legacy projects where, most likely, no specification or design models are available but rather only a few drawings as part of the system documentation.
@@
Additionally, in a low-code context, which focuses on reducing the amount of hand-coding required to create applications [10], these models can be further processed to (semi)automatically generate software. Following these footsteps, the term "low-modeling" [11] was also created to describe techniques and tools that accelerate the modeling process, effectively speeding up the low-code pipeline.
We believe that a tool capable of converting given images to UML diagrams could speed up this tedious process and quickly provide initial models that could be extended or further used, effectively enabling the low-modeling of software.
@@
In this paper, we explore the capabilities of LLMs with image processing capabilities to act as image to UML converters. To be precise, we conducted experiments on various images of drawn UML diagrams using different visual LLMs and evaluate the correctness and completeness of the generated results. We proceeded to discuss and interpret the results based on the blackbox nature of the LLMs. Our findings reveal that GPT4-V provided the best results in terms of correctness and completeness, and indeed that the usage of LLMs to convert images to UML provides positive results, although with the need of keeping the human in the loop and dependend on the used LLM.
@@ 2 State of the art
Other works abstract the task further by either only providing a set of information from given diagrams [14, 15], such as the location and text inside classes, or only attempt to classify the type of UML diagram and not the content itself using deep learning techniques [16, 17]. Not a lot of research focuses on the ability to extract information from images of hand-drawn UML diagrams and existing attempts are outdated and do not hold up to the constant technological progress or focus on drawings made with a specific tool and not on actual paper [18, 19].
As previously mentioned, attempts at generating UML models using LLMs have already been made [4, 5], both using the PlantUML notation as output. Yet, these two attempts only tackle textual input. To the best of our knowledge, we are the first to leverage LLMs’ image recognition capabilities to generate UML models.
@@ 3 Research questions
Our main goal is to evaluate vision LLMs’ capabilities to process and transform images of UML diagrams into a computer readable format while also exploring possible variables that affect the transformation. We focus on hand-drawn UML diagrams as we assume that the transformation of images of hand-drawn UML diagrams into a computer readable format is more difficult than the same task for images of UML diagrams that were created with CASE tools, as drawings generally contain more inconsistencies in terms of handwriting, lines, etc. This assumption implies that any results obtained with images of hand-drawn UML diagrams can be used as an approximate result for images of UML diagrams created with CASE tools. Moreover, we will primarily focus on UML class diagrams, as these seem to be the most popular type of UML diagram [20, 21].
We partially inspire ourselves from the experiment conducted in [5] and formulate the following Research Questions (RQ):
@@ 4.1 Setup
To answer the RQs, we iterate through different examples of UML class diagrams and evaluate the produced results. For that purpose, we defined 4 diagrams, where the first 3 are denoted by a steady increase in difficulty due to the addition of elements and concepts, and the final diagram represents a UML class diagram with a correct syntax but that represents a model that, semantically, is not representing a realistic domain. These hand-drawn images are fed into an LLM with a corresponding prompt requesting a generation of the understood model using a concrete notation as output. As concrete notation, we opted for the PlantUML [^4] https://plantuml.com/ notation, as it is a text-based diagramming tool that enables the creation of UML diagrams such as class diagrams using a simple and intuitive syntax. It is useful for quickly and efficiently visualizing and communicating the structure of software systems and can be ingested by generators to produce applications.
@@
We also decided to define multiple prompts with each prompt increasing in the level of detail of the provided. In increasing order, the used prompts are:
@@
Preliminary tests showed that while LLMs provide nondeterministic results when generating the PlantUML class diagrams, the degree of variance between each attempt for a given example and LLM was not large. Therefore, we decided that per prompt and per LLM, evaluating three runs should be enough to give an adequate performance overview. Each attempt will be done in an empty conversation to avoid any kind of influence from previous messages.
@@ 4.2 Results
Furthermore, Table 2 summarizes the number of times wrong PlantUML syntax was generated and the times the LLM refused to generate PlantUML code. CogVLM produced the highest number of outputs with syntax errors, closely followed by Gemini Pro and Ultra. On the other hand, GPT-4V produced not a single syntax error after 36 generations.
Regarding the nature of the syntax errors, the Gemini models often had trouble with the notation to define inter-class relationships. An example would be for inheritance, instead of the correct syntax "class Fish extends Animal{…}" Gemini Pro would generate "class Fish inherits Animal{…}" or "class Fish <|– Animal{…}". As for CogVLM, simple class definitions such as "class Duck{}" would already cause problems as it would attempt to define classes as "Duck{}".
Additionally, only Gemini models refused to generate solutions for some of the images. Especially level 4 resulted in the most outputs without generation. Notably, Gemini models deemed the given syntactically correct but semantically questionable input as not being real UML models and therfore refusing to process them. An example of such a response can be seen in Figure 2. The nonsensical diagram also seems to have caused a high number of syntax errors for the Gemini Pro and CogVLM model.
@@
RQ1: Are LLMs capable of providing a complete (classes, relationships, textual content,…) re-creation of a given UML diagram?
When generated without syntax errors, LLMs managed to re-create the given UML class diagram with a variable degree of accuracy, depending on the used LLM. While sometimes some elements were either missing, wrongly interpreted or hallucinated, the resulting output of the best result ended up providing a faithful recreation of the given input. Nevertheless, there is a huge variation in the quality of the results so the choice of the LLM is key.
RQ2: Do LLMs respect the syntax of the chosen notation for the output?
As we saw, GPT-4V flawlessly provided PlantUML code containing no syntax errors. The same cannot be said about the other tested models that tended to provide wrong syntax for more around 50% of cases. Overall, LLMs are able to provide correct PlantUML syntax, it just depends which one is used.
RQ3: Does complexity of the given diagram affect the results?
As highlighted, the results present in Table 1 show that the complexity of the input diagram, defined by a higher number of elements and concepts, negatively affect the accuracy of the generated result. This correlation seems to count for every LLM used in the experiment.
While GPT-4V does not seem to be affected by the semantic correctness of the given input, the other LLMs are strongly affected by the semantic correctness or logic behind the given input.
RQ5: Does descriptiveness of the prompt affect the results?
As stated, our results show that the least descriptive prompt led to the best results overall and the most descriptive prompt to the worst. While counterintuitive, as prompt engineering usually advocates that more descriptive prompts lead to better results [1], based on our results, we would argue that the descriptiveness negatively affects the results for the PlantUML generation task.
@@ 5 Discussion
Inconsistency in the output syntax
Gemini is inconsistent with the quality of the output syntax. While sometimes it is able to generate the proper one (showing that in fact it knows it), it also often generates wrong syntax.
We can force it to correct the syntax by explicitly showing it the right syntax to use. But just telling Gemini that the syntax is wrong does not fix the issue (even if we know that in fact it is able to generate proper syntax).
This showcases the importance of prompt engineering and that different LLMs expect different prompting strategies to complete (and correct) tasks successfully. At the same time, one could use this as a metric to rate the quality of LLMs in future experiments, as a LLM completing a task without needing a reminder of the syntax of a notation is less tedious to use.
Imposed view of "real" UML class diagrams
The Gemini models sometimes refused to generate any PlantUML code as it deemed the given input as not being real UML. This phenomena mostly occurred when attempting to convert the "nonsensical" UML class diagram. An example of such a response can be seen in Figure 2. Interestingly, appending the sentence "Ignore the semantics." to the prompt resulted in actual PlantUML generation taking place. This again highlights the importance of prompt engineering, but also hints at default restrictions being implemented into Gemini, imposing its truth onto the user except when specified otherwise.
@@
Choosing an output notation with enough training data available
@@
Beside PlantUML, other textual notations, such as Umple or yuml, exist to define UML diagrams. Yet, informal testing has shown that the tested LLMs seemed to have the least trouble with PlantUML. While no thorough systematic tests were done on the different notations, we still see that the choice of the notation is important, as it affects the quality of the output. Furthermore, the better performance when using the PlantUML notation seems to correlate with the popularity of the notation, as PlantUML is considered one of the most popular UML tools available [^5] https://modeling-languages.com/text-uml-tools-complete-list/. Arguably, this leads to more PlantUML data being available and thus, more PlantUML data contained in the LLMs’ training data.
Keeping the human in the loop
@@
Overall, while the Gemini models and CogVLM manage to provide a sort of base prototype for the given images, a lot of syntax and general adjustments are needed as compared to GPT-4V that usually provides an almost correct end-result in one go, without needing to complicate the prompt. We would argue that GPT-4V is the most suitable for the given task. Nevertheless, the minor errors still require human correction, thus, following the human-in-the-loop paradigm is currently still necessary.
Figure 2: Example of Gemini Ultra refusing to generate PlantUML code
5.1 Threats to Validity
We will now briefly discuss aspects that might affect this experiment’s validity.
5.1.1 Threats to internal validity
Chosen prompts
The chosen prompts might not have been the best ones for such a task, as is evident by Gemini’s change in response to the given nonsensical UML class diagram when appending "Ignore the semantics." to the prompt. Another experiment could aim to determine the best prompt to create UML class diagrams out of images depending on the used LLM.
Grading scheme
@@ 5.1.1 Threats to internal validity
The grading scheme used to evaluate the output potentially may not completely reflect the performance of the LLMs. Aspects such as every mistake, excluding syntax errors, being worth 1 mistake could be seen as limited. Some might argue that forgetting an entire class might be worth more than just forgetting to add an attribute. Additionally, when syntax errors occurred, we marked the result of the attempt as an error. This resulted in outputs with correct syntax but with a lot of mistakes and hallucinations to be evaluated as a better result than those with syntax errors. One could question whether it is fair to say that something that does not compile but has less mistakes is worse than something that does compile but contains a lot of unwanted and wrong elements. In this case, we would argue that sometimes, creating something from scratch might prove easier than needing to change a lot of mistakes in a given prototype.
Scalability to larger diagrams
@@
Finally, the used UML class diagrams were fairly small in nature and do not cover more complex software systems. Our observed results might not be applicable to larger UML class diagrams. We would still argue that the experiment itself is realistic in the sense that, in collaborative sessions, small or very abstract prototypes, or small components of a larger system are sketched and not big and complex UML class diagrams.
5.1.2 Threats to external validity
Nondeterministic nature of LLMs and sample size

### v2

# First Large Language Models for Model-Driven Engineering Workshop (LLM4MDE 2024), Enschede, Netherlands

@@ Abstract
In software engineering processes, systems are first specified using a modeling language such as UML. These initial designs are often collaboratively created, many times in meetings where different domain experts use whiteboards, paper or other types of quick supports to create drawings and blueprints that then will need to be formalized. These proper, machine-readable, models are key to ensure models can be part of automated processes (e.g. input of a low-code generation pipeline, a model-based testing system, …). But going from hand-drawn diagrams to actual models is a time-consuming process that sometimes ends up with such drawings just added as informal images to the software documentation, reducing their value a lot. To avoid this tedious task, we explore the usage of Large Language Models (LLM) to generate the formal representation of (UML) models from a given drawing. More specifically, we have evaluated the capabilities of different LLMs to convert images of (hand-drawn) UML class diagrams into the actual models represented in the images. While the results are good enough to use such an approach as part of a model-driven engineering pipeline we also highlight some of their current limitations and the need to keep the human in the loop to overcome those limitations.
keywords:
[…]
@@ 1 Introduction
The continuous progress and rise of Large Language Models (LLM) has led to their integration into various tasks in various domains, such as medical advice consultation in healthcare, writing or reading assistance in education, legal interpretation and reasoning in Law or financial reasoning [1]. Computer science is not an exception, with various LLMs being used in software development as programming assistants [2], but also in software modeling as a model creation tool from natural language [3, 4], showing promising results on both fronts. Beyond textual input, some LLMs additionally provide support for image input, also described as visual LLMs. Generally, OpenAI’s GPT-4V [^1] https://openai.com/research/gpt-4v-system-card and Google’s Gemini (Pro/Ultra) [^2] https://deepmind.google/technologies/gemini are seen as the best publicly available multimodal LLM. Unfortunately, these are either hidden behind a paywall or only accessible through specific platforms’ interfaces. CogVLM [5] acts as an open-source alternative, providing not only its source code but also a free chat interface to interact with the model. These multi-modal LLMs further broadened the range of possible applications, such as tools like Design2Code [6] or Make-real [^3] https://github.com/tldraw/make-real that allow for the generation of HTML, CSS and JS code from either screenshots of web pages or mock-ups. In the domain of software modeling, we believe that these image reading capabilities could also be leveraged to facilitate and accelerate the modeling process itself.
@@
Indeed, one of the biggest hurdles of modeling tools is the usability of such tools [7]. Take the example of UML class diagrams that are used to represent the structure of software. These are usually collaboratively sketched on a whiteboard in an attempt to crystallize the initial design idea of an application. Yet, once drawn, there still is the need to transform the drawing into a more polished form, for example for documentation purposes. Additionally, one might also want a computer readable format to use the UML class diagram for code generation. The same goes for migrating legacy projects where, most likely, no specification is available but rather only a few drawings as part of the system documentation.
@@
In a low-code context, which focuses on reducing the amount of hand-coding required to create applications, these models can be further processed to (semi)automatically generate software. Following these footsteps, the term "low-modeling" [8] was also created to describe techniques and tools that accelerate the modeling process, effectively speeding up the low-code pipeline. We believe that a tool capable of converting given images to UML diagrams could speed up this tedious process and quickly provide initial models that could be extended or further used, effectively enabling the low-modeling of software.
@@
In this paper, we explore the capabilities of LLMs with image processing capabilities to act as image to UML converters, contributing towards the hybridization between Software Engineering and LLMs [9]. We conducted experiments on various images of drawn UML diagrams using visual LLMs and evaluate the correctness and completeness of the generated results. Our findings reveal that GPT4-V provided the best results in terms of correctness and completeness, and indeed that the usage of LLMs to convert images to UML provides positive results, although with the need of keeping the human in the loop and dependent on the used LLM.
@@ 2 State of the art
Other works abstract the task further by either only providing a set of information from given diagrams [12], such as the location and text inside classes, or only attempt to classify the type of UML diagram and not the content itself using deep learning techniques [13]. Not a lot of research focuses on the ability to extract information from images of hand-drawn UML diagrams and existing attempts are outdated and do not hold up to the constant technological progress or focus on drawings made with a specific tool and not on actual paper [14, 15].
As previously mentioned, attempts at generating UML models using LLMs have already been made [3, 4] using the PlantUML notation as output, yet, these two attempts only tackle textual input. To the best of our knowledge, we are the first to leverage LLMs’ image recognition capabilities to generate UML models.
@@ 3 Research questions
Our main goal is to evaluate vision LLMs’ capabilities to process and transform images of UML diagrams into a computer readable format while also exploring possible variables that affect the transformation. We focus on hand-drawn UML diagrams as we assume that the transformation of images of hand-drawn UML diagrams into a computer readable format is more difficult than the same task for images of UML diagrams that were created with CASE tools, as drawings generally contain more inconsistencies in terms of handwriting, lines, etc. This assumption implies that any results obtained with images of hand-drawn UML diagrams can be used as an approximate result for images of UML diagrams created with CASE tools. Moreover, we will primarily focus on UML class diagrams, as these seem to be the most popular type of UML diagram [16]. We partially inspire ourselves from the experiment conducted in [4] and formulate the following Research Questions (RQ):
@@ 4.1 Setup
To answer the RQs, we iterate through different examples of UML class diagrams and evaluate the produced results. For that purpose, we defined 4 diagrams, where the first 3 are denoted by a steady increase in difficulty due to the addition of elements and concepts, and the final diagram represents a UML class diagram with a correct syntax but that represents a model that, semantically, is not representing a realistic domain. These hand-drawn images, all created by the same person, are fed into an LLM with a corresponding prompt requesting a generation of the understood model using a concrete notation as output. We opted for the aforementioned LLMs GTP-4V, Gemini (Pro and Ultra) and CogVLM with its default configuration (top_p=0.40, temperature=0.80, top_k=1). As concrete notation, we opted for the PlantUML [^4] https://plantuml.com/ notation, as it is a text-based diagramming tool that enables the creation of UML diagrams such as class diagrams using a simple and intuitive syntax, ingestable by generators to produce applications.
@@
We also crafted several prompts, each offering increasing levels of detail to describe the task. In increasing order, the used prompts are:
@@
Preliminary tests showed that while LLMs provide nondeterministic results when generating the PlantUML class diagrams, the degree of variance between each attempt for a given example and LLM was not large. Therefore, we decided that per prompt and per LLM, evaluating three runs should be enough to give an adequate performance overview. Each attempt has been done in an empty conversation to avoid any kind of influence from previous messages.
@@ 4.2 Results
Furthermore, Table 2 summarizes the number of times wrong PlantUML syntax was generated and the times the LLM refused to generate PlantUML code. CogVLM produced the highest number of outputs with syntax errors, closely followed by Gemini Pro and Ultra. On the other hand, GPT-4V produced not a single syntax error after 36 generations. Regarding the nature of the syntax errors, the Gemini models often had trouble with the notation to define inter-class relationships. An example would be for inheritance, instead of the correct syntax "class Fish extends Animal{…}" Gemini Pro would generate "class Fish <|– Animal{…}". As for CogVLM, simple class definitions such as "class Duck{}" would already cause problems as it would attempt to define classes as "Duck{}".
Additionally, only Gemini models refused to generate solutions for some of the images. Especially level 4 resulted in the most outputs without generation. Notably, Gemini models deemed the given syntactically correct but semantically questionable input as not being real UML models and therefore refusing to process them. An example of such a response would be "Unfortunately, I cannot translate the hand-drawn UML class diagram into PlantUML notation based on the image you provided. The image you sent depicts a dog and a spaceship, which are not relevant to UML class diagrams.". The nonsensical diagram also seems to have caused a high number of syntax errors for the Gemini Pro and CogVLM model.
@@
RQ1: Are LLMs capable of providing a complete (classes, relationships, textual content,…) re-creation of a given UML diagram? When generated without syntax errors, LLMs managed to re-create the given UML class diagram with a variable degree of accuracy, depending on the used LLM. While sometimes some elements were either missing, wrongly interpreted or hallucinated, the resulting output of the best result ended up providing a faithful recreation of the given input. Nevertheless, there is a huge variation in the quality of the results so the choice of the LLM is key.
RQ2: Do LLMs respect the syntax of the chosen notation for the output? As we saw, GPT-4V flawlessly provided PlantUML code containing no syntax errors. The same cannot be said about the other tested models that tended to provide wrong syntax for more around 50% of cases. Overall, LLMs are able to provide correct PlantUML syntax, depending on the LLM.
RQ3: Does complexity of the given diagram affect the results? As highlighted, the results present in Table 1 show that the complexity of the input diagram, defined by a higher number of elements and concepts, negatively affect the accuracy of the generated result. This correlation seems to count for every LLM used in the experiment.
RQ4: Does semantic correctness of the given diagram affect the results? While GPT-4V does not seem to be affected by the semantic correctness of the given input, the other LLMs are strongly affected by the semantic correctness or logic behind the given input.
RQ5: Does descriptiveness of the prompt affect the results? As stated, our results show that the least descriptive prompt led to the best results overall and the most descriptive prompt to the worst. While counterintuitive, as prompt engineering usually advocates that more descriptive prompts lead to better results [1], based on our results, we would argue that the descriptiveness negatively affects the results for the PlantUML generation task.
@@ 5 Discussion
Inconsistency in the output syntax. Gemini is inconsistent with the quality of the output syntax. While sometimes it is able to generate the proper one (showing that in fact it knows it), it also often generates wrong syntax. We can force it to correct the syntax by explicitly showing it the right syntax to use, but just telling Gemini that the syntax is wrong does not fix the issue (even if we know that in fact it is able to generate proper syntax). This showcases the importance of prompt engineering and that different LLMs expect different prompting strategies to complete (and correct) tasks successfully. At the same time, one could use this as a metric to rate the quality of LLMs in future experiments, as a LLM completing a task without needing a reminder of the syntax of a notation is less tedious to use.
Imposed view of "real" UML class diagrams. The Gemini models sometimes refused to generate any PlantUML code as it deemed the given input as not being real UML. This phenomena mostly occurred when attempting to convert the "nonsensical" UML class diagram. Interestingly, appending the sentence "Ignore the semantics." to the prompt resulted in actual PlantUML generation taking place. This again highlights the importance of prompt engineering, but also hints at default restrictions being implemented into Gemini, imposing its truth onto the user except when specified otherwise.
@@
Choosing an output notation with enough training data available. Beside PlantUML, other textual notations, such as Umple or yuml, exist to define UML diagrams. Yet, informal testing has shown that the tested LLMs seemed to have the least trouble with PlantUML. While no thorough systematic tests were done on the different notations, we still see that the choice of the notation is important, as it affects the quality of the output. Furthermore, the better performance when using the PlantUML notation seems to correlate with the popularity of the notation, as PlantUML is considered one of the most popular UML tools available [^5] https://modeling-languages.com/text-uml-tools-complete-list/. Arguably, this leads to more PlantUML data being available and thus, more PlantUML data contained in the LLMs’ training data.
@@
Keeping the human in the loop. Overall, while the Gemini models and CogVLM manage to provide a sort of base prototype for the given images, a lot of syntax and general adjustments are needed as compared to GPT-4V that usually provides an almost correct end-result in one go, without needing to complicate the prompt. We would argue that GPT-4V is the most suitable for the given task. Nevertheless, the minor errors still require human correction, thus, following the human-in-the-loop paradigm is currently still necessary.
5.1 Threats to internal validity
@@
Grading scheme. The grading scheme used to evaluate the output potentially may not completely reflect the performance of the LLMs. Aspects such as every mistake, excluding syntax errors, being worth 1 mistake could be seen as limited. Some might argue that forgetting an entire class might be worth more than just forgetting to add an attribute. Additionally, when syntax errors occurred, we marked the result of the attempt as an error. This resulted in outputs with correct syntax but with a lot of mistakes and hallucinations to be evaluated as a better result than those with syntax errors. One could question whether it is fair to say that something that does not compile but has less mistakes is worse than something that does compile but contains a lot of unwanted and wrong elements. In this case, we would argue that sometimes, creating something from scratch might prove easier than needing to change a lot of mistakes in a given prototype.
@@
Scalability to larger diagrams. Finally, the used UML class diagrams were fairly small in nature and do not cover more complex software systems. Our observed results might not be applicable to larger UML class diagrams. We would still argue that the experiment itself is realistic in the sense that, in collaborative sessions, small or very abstract prototypes, or small components of a larger system are sketched and not big and complex UML class diagrams.
Chosen prompts. The chosen prompts might not have been the best ones for such a task, as is evident by Gemini’s change in response to the given nonsensical UML class diagram when appending "Ignore the semantics." to the prompt. Another experiment could aim to determine the best prompt to create UML class diagrams out of images depending on the used LLM.
5.2 Threats to external validity